OpenStudy (anonymous):

Which set of equations would be used to solve this word problem? An airplane flew 5 hours with a 50 mph head wind. The return trip with a tail wind of the same speed took 3 hours. Find the speed of the plane in still air. Answer None of these systems can be used to solve this problem. d = 5(r + 50) and d = 3(r - 50) d = 8(r + 50) d = 5(r - 50) and d = 3(r + 50)

OpenStudy (anonymous):

d = 5(r+50) and d=3(r-50)